# Efficient Analysis of Latent Spaces in Heterogeneous Networks

### v2

@@ Abstract
This work proposes a unified framework for efficient estimation under latent space modeling of heterogeneous networks. We consider a class of latent space models that decompose latent vectors into shared and network-specific components across networks. We develop a novel procedure that first identifies the shared latent vectors and further refines estimates through efficient score equations to achieve statistical efficiency. Oracle error rates for estimating the shared and heterogeneous latent vectors are established simultaneously. The analysis framework offers remarkable flexibility, accommodating various types of edge weights under exponential family distributions.
@@ 1 Introduction
In recent years, analyses of multiview or multiplex networks (Kivelä et al.,, 2014) open up new opportunities for understanding complex systems across diverse scientific endeavors (Salter-Townshend and McCormick,, 2017), including social science (Lazega,, 2001; Banerjee et al.,, 2013), microbiome (Gould et al.,, 2018), and neuroscience (Wen et al.,, 2022). In these applications, it is common to observe multiple networks sharing a common set of nodes, with each edge representing specific types of connections or relationships between node pairs. Studying multiple networks holistically enables a unified understanding of multifaceted and interconnected systems.
@@
Latent space modeling has emerged as an important approach in network analysis (Hoff et al.,, 2002; Matias and Robin,, 2014). In this framework, each node is mapped to a vector in a latent space, and the connectivity strength between two nodes is measured through a function between two corresponding latent vectors. While latent spaces can be discrete, they are more commonly referred to as stochastic block models and exhibit unique properties from discreteness (Holland et al.,, 1983). Throughout this paper, we focus on continuous latent vectors and consider latent spaces as Euclidean spaces following the convention in the literature (Hoff et al.,, 2002).
@@
For multiple or multilayer networks, various developments under latent space modeling have been proposed from diverse perspectives. Bayesian approaches treat latent vectors as random variables (Hoff,, 2011; Salter-Townshend and McCormick,, 2017), while frequentist approaches view them as fixed parameters (Arroyo et al.,, 2021; Jones and Rubin-Delanchy,, 2020; Nielsen and Witten,, 2018; Zheng and Tang,, 2022; Zhang et al.,, 2020; MacDonald et al.,, 2022). Our work aligns with the latter, focusing on estimating the fixed latent vectors.
@@
Existing studies have consistently shown that multiple networks often share underlying structures while simultaneously exhibiting significant network-specific heterogeneity (Wang et al.,, 2019; Zhang et al.,, 2020; Arroyo et al.,, 2021; Chen et al.,, 2022; He et al.,, 2023; MacDonald et al.,, 2022; Lei and Lin,, 2023). When analyzing a collection of heterogeneous networks, leveraging their shared information is crucial for enhancing statistical efficiency, whereas accurately characterizing unique features is equally important. The coexistence of shared and unique patterns calls for analysis capable of disentangling the intertwined structural features and still fully exploiting statistical efficiency. A recent advancement by MacDonald et al., (2022) introduced a general class of latent space models for multiple networks that decompose latent vectors into shared and network-specific components. Despite flexibility of the models, fundamental statistical limits under them remain unclear. Specifically, a critical statistical question for jointly analyzing multiple networks is: Can pooling heterogeneous networks improve the efficiency of estimating latent vectors, and if so, how can such gains be effectively achieved?
@@
To address this, we develop a unified framework of efficient analysis under the latent space modeling of multiple heterogeneous networks. We propose a novel strategy that first hunts for the shared latent space across networks and then applies a refinement to achieve oracle efficiency. In this way, challenges from identifying shared space and achieving efficiency can be tackled separately, enhancing modeling and analytical flexibility. To theoretically quantify the efficiency gain, we develop novel techniques that disentangle the distinct oracle estimation error rates of shared and network-specific spaces through analyzing complex efficient score equations. Our proposed framework is versatile and encompasses a broad range of edge weight distributions within exponential family models.
@@
We summarize notations that will be used throughout this paper. Given two sequences of real numbers $\{g_{n}\}$ and $\{h_{n}\}$, $g_{n}=O(h_{n})$ and $g_{n}\lesssim h_{n}$ represent that $|g_{n}|\leqslant c_{1}|h_{n}|$ for a constant $c_{1}>0$, $g_{n}\ll h_{n}$ and $g_{n}=o(h_{n})$ mean $\lim_{n\to\infty}g_{n}/h_{n}=0$, and $g_{n}\gg h_{n}$ means $\lim_{n\to\infty}h_{n}/g_{n}=0$. Consider a sequence of random variables $\{X_{n}\}$. Write $X_{n}=O_{p}(g_{n})$ if for any $\epsilon>0$, there exists finite $M>0$ such that $\sup_{n}\Pr(|X_{n}/g_{n}|>M)<\epsilon$; write $X_{n}=o_{p}(g_{n})$ if for any $\epsilon>0$, $\lim_{n\to\infty}\Pr(|X_{n}/g_{n}|>\epsilon)=0$. For a vector $x=(x_{i})_{i=1}^{n}\in\mathbb{R}^{n}$, let $\|x\|_{2}=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$ and $\|x\|_{\infty}=\max_{1\leqslant i\leqslant n}|x_{i}|$. For a matrix $X=(x_{ij})_{1\leqslant i\leqslant n,1\leqslant j\leqslant m}\in\mathbb{R}^{n\times m}$, let $\|X\|_{\mathrm{F}}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}x_{ij}^{2}}$, $\|X\|_{\operatorname{op}}=\sup_{\|v\|_{2}=1}\|Xv\|_{2}$, and $\|X\|_{2\to\infty}=\sup_{\|v\|_{2}=1}\|Xv\|_{\infty}$.
@@ 2.1 Latent space model with shared factors
This paper focuses on multiple networks where each network is observed on a common set of $n$ codes without inter-layer connections. Suppose we observe $T$ undirected networks on a common set of $n$ nodes, represented by $n\times n$ adjacency matrices $\{\mathbf{A}_{t}=(A_{t,ij})_{n\times n}:t=1,\ldots,T\}$. For weighted networks, $A_{t,ij}$ is the weight value, which could be continuous or a count number. For unweighted networks, $A_{t,ij}$ is binary indicating that the connection between nodes $i$ and $j$ exists or not.
@@
To model interactions between nodes, we adopt the popular latent space modeling framework (Hoff et al.,, 2002). Assume there exist latent vectors $y_{t,i}\in\mathbb{R}^{d_{t}}$ for $i=1,\ldots,n$ and $t=1,\ldots,T$ such that conditioning on $y_{t,i}$’s, network edges $A_{t,ij}$ are independently generated. To flexibly model different types of weight values, we allow $A_{t,ij}$ to be generated from a natural exponential family of distributions. In particular, let $\mathcal{F}_{\nu}(\,\cdot\,;\,\theta\,)$ represent the probability density function of a one-parameter exponential family distribution with the natural parameter $\theta$ and the log-partition function $\nu(\cdot)$, i.e., $\mathcal{F}_{\nu}(x;\theta)\propto\exp\{\theta x-\nu(\theta)\}$. We consider
[…]
independently, where $\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle$ models the interaction effect between nodes $i$ and $j$ similarly to the popular low-rank models in the literature (Hoff et al.,, 2002; Ma et al.,, 2020).
@@
For each node $i$, we assume that the latent vectors $\{y_{t,i}:t=1,\ldots,T\}$ may include shared and heterogeneous components over $T$ networks following MacDonald et al., (2022). Specifically, we can write $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}\in\mathbb{R}^{k+k_{t}}$, where $k+k_{t}=d_{t}$, $z_{i}\in\mathbb{R}^{k}$ represents a shared component over $T$ networks and does not change with respect to the index $t$, and $w_{t,i}\in\mathbb{R}^{k_{t}}$ represents a heterogeneous component that may differ over $1\leqslant t\leqslant T$. Then
$\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle=\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle.$ (2)
Model (1) assumes a common type of distribution function $\mathcal{F}$, which is used for notational simplicity. It is worth mentioning that the proposed analysis framework can be readily generalized when networks follow different types of distributions.
@@
Let $\Theta_{t}=(\Theta_{t,ij})_{1\leqslant i,j\leqslant n}\in\mathbb{R}^{n\times n}$ denote the matrix of the natural parameters for $t$-th network. Then (1) and (2) lead to an equivalent low-rank matrix representation
$g\{\mathbb{E}(\mathbf{A}_{t}\mid Y_{t})\}=\Theta_{t}=Y_{t}Y_{t}^{\top}=ZZ^{\top}+W_{t}W_{t}^{\top},\hskip 15.00002pt1\leqslant t\leqslant T,$ (3)
where $g(\cdot)=(\nu^{\prime})^{-1}(\cdot)$ represents the link function of the given exponential family distribution $\mathcal{F}_{\nu}(\cdot)$ in (1), operating entrywisely on a matrix argument, and we define $Y_{t}=[Z,\,W_{t}]\in\mathbb{R}^{n\times d_{t}},$ $Z=\left(z_{1},\ldots,z_{n}\right)^{\top}\in\mathbb{R}^{n\times k},$ and $W_{t}=\left(w_{t,1},\ldots,w_{t,n}\right)^{\top}\in\mathbb{R}^{n\times k_{t}}.$
@@ 2.2 Identifiability
By (3), the parameters $Z$ and $W_{t}$ can only be identified up to orthogonal transformation, i.e., $ZZ^{\top}=ZQQ^{\top}Z^{\top}$ and $W_{t}W_{t}^{\top}=W_{t}Q_{t}Q_{t}^{\top}W_{t}^{\top}$ for any $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$, where $\mathcal{O}(k)=\left\{Q\in\mathbb{R}^{k\times k}:QQ^{\top}=\mathrm{I}_{k}\right\}$ represents the set of $k\times k$ orthogonal matrices, and $\mathrm{I}_{k}$ denotes $k\times k$ identity matrix. We say that parameters $Z$ and $W_{t}$ in the model (3) are identifiable up to orthogonal group transformation if, for any two sets of parameters $\{Z,W_{1},\ldots,W_{T}\}$ and $\{Z^{\prime},W_{1}^{\prime},\ldots,W_{T}^{\prime}\}$ yielding the same model (3), there exist $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$ such that $Z=Z^{\prime}Q$ and $W_{t}=W^{\prime}_{t}Q_{t}$ for $1\leqslant t\leqslant T$. We next establish a sufficient condition for the identifiability of the model (3).
@@
    there exist $1\leqslant t<s\leqslant T$ such that the columns of $[Z,W_{t},W_{s}]$ are linearly independent.
@@
Proposition 1 is adapted from MacDonald et al., (2022) but relaxes the connectivity assumption of the inter-layer graph under the model (2). Given (i) and (ii) in Proposition 1, the column span of $Z$ equals the intersection of all column spans of $Y_{t}$ across $1\leqslant t\leqslant T$.
@@ 2.3 Efficient analysis of latent spaces and challenges
For estimating the shared factor $Z$ under (3), we expect a similar semiparametric oracle error rate can be established. To illustrate the intuition, we consider an oracle scenario when heterogeneous (nuisance) parameters $w_{t,i}$’s are known. In this case, each shared (target) parameter $z_{i}$ could be measured by $nT$ independent edges $\{A_{t,ij}:j=1,\ldots,n;t=1,\ldots,T\}$. The oracle squared estimation error rate of each latent vector $z_{i}$ can be expected to be of the order of $O_{p}(k/(nT))$, the ratio between the parameter number and the effective sample size (Portnoy,, 1988). Thus the aggregated oracle estimation error of $n$ latent vectors $Z$ is expected to be $O_{p}(n\times k/(nT))=O_{p}(1/T)$ under fixed $k$. To establish the above oracle error rate under (3), technical challenges similar to those in He et al., (2023) persist.
@@
    Oracle error rates for estimating target and nuisance parameters are different, but their errors are entangled in the analysis. Under (2), each nuisance $w_{t,i}$ is measured by $n$ independent edges: $\left\{A_{t,ij}:j=1,\ldots,n\right\}$, so the oracle squared error rate of each $w_{t,i}$ is expected to be $O_{p}(1/n)$, larger than $O_{p}(1/(nT))$ of each $z_{i}$. In theoretical analysis, it is difficult to separate the error of $z_{i}$ from that of $w_{t,i}$ due to several reasons. One is that the individual components $w_{t,i}$’s encode both node-specific heterogeneity over $i$ and network-specific heterogeneity over $t$. As a result, classical partial likelihood approach (Andersen and Gill,, 1982) cannot directly remove the heterogeneous factors. Moreover, when the link function in (3) is non-linear, such as the logistic or exponential function, semiparametric oracle rates for $Z$ cannot be easily obtained by spectral-based analyses (Arroyo et al.,, 2021) or analysis under Gaussian distribution (MacDonald et al.,, 2022).
@@
    Target parameters are unidentifiable. As argued in Section 2.2, the shared (target) latent space $Z$ is unidentifiable in the Euclidean space. Actually, the intrinsic space of $Z$ is a quotient set with the equivalence relation up to the orthogonal group transformation. To evaluate the discrepancy between a target matrix $Z^{\star}$ and an estimate $\hat{Z}$, the Euclidean distance between two matrices is inappropriate, and a natural distance is $\mathrm{dist}(\hat{Z},Z^{\star})=\min_{Q\in\mathcal{O}(k)}\|\hat{Z}-Z^{\star}Q\|_{\mathrm{F}}$.
@@
    Both target and nuisance parameters are high-dimensional. Specifically, under (3), the dimension of $Z$ increases with respect to the network size $n$, and the total number of parameters in $[W_{1},\ldots,W_{T}]$ grows with respect to both $n$ and $T$.
On the other hand, our model (3) introduces unique challenges and properties, making existing analyses inapplicable and necessitating new developments of methodology and theory. We summarize critical distinctions into four key aspects.
@@
    Intricate factor interrelationships. The relationship between the shared (target) space $Z$ and individual (nuisance) spaces $W_{t}$’s exhibits unprecedented complexity. In He et al., (2023), the target and nuisance parameters are latent factors and baseline degrees, respectively, which characterize intrinsically orthogonal information that can enhance theoretical analysis. In contrast, our model (3) confronts a far more intricate landscape where factor interrelationships defy simple characterizations, which will be further elaborated in Remark 3. This complexity greatly amplifies the challenge of separating estimation errors across different $1+T$ factor groups.

### v3

@@ Abstract
This work proposes a unified framework for efficient estimation under latent space modeling of heterogeneous networks. We consider a class of latent space models that decompose latent vectors into shared and network-specific components across networks. We develop a novel procedure that first identifies the shared latent vectors and further refines estimates through efficient score equations to achieve statistical efficiency. Oracle error rates for estimating the shared and heterogeneous latent vectors are established simultaneously. The analysis framework offers remarkable flexibility, accommodating various types of edge weights under general distributions.
@@ 1 Introduction
In recent years, analyses of multiview or multiplex networks (Kivelä et al.,, 2014) open up new opportunities for understanding complex systems across diverse scientific endeavors (Salter-Townshend and McCormick,, 2017), including social science (Lazega,, 2001; Banerjee et al.,, 2013), microbiome (Gould et al.,, 2018), and neuroscience (Wen et al.,, 2022). In these applications, it is common to observe multiple networks sharing a common set of nodes, with each edge representing specific types of relationships between node pairs. Studying multiple networks enables a unified understanding of multifaceted and interconnected systems.
@@
Latent space modeling has emerged as an important approach in network analysis (Hoff et al.,, 2002; Matias and Robin,, 2014). In this framework, each node is mapped to a vector in a latent space, and the connectivity strength between two nodes is parametrized through a function between two corresponding latent vectors. While latent spaces can be discrete, they are more commonly referred to as stochastic block models and exhibit unique properties from discreteness (Holland et al.,, 1983). Throughout this paper, we focus on continuous latent vectors and consider latent spaces as Euclidean spaces following the convention in the literature (Hoff et al.,, 2002).
@@
For multiple or multilayer networks, various developments under latent space modeling have been proposed, including Bayesian approaches (Hoff,, 2011; Salter-Townshend and McCormick,, 2017), and frequentist approaches (Arroyo et al.,, 2021; Jones and Rubin-Delanchy,, 2020; Nielsen and Witten,, 2018; Zheng and Tang,, 2022; Zhang et al.,, 2020; MacDonald et al.,, 2022). Our work aligns with the latter, focusing on estimating fixed latent vectors.
@@
Existing studies have consistently shown that multiple networks often share underlying structures while simultaneously exhibiting significant network-specific heterogeneity (Wang et al.,, 2019; Zhang et al.,, 2020; Arroyo et al.,, 2021; Chen et al.,, 2022; He et al.,, 2025; MacDonald et al.,, 2022). When analyzing a collection of heterogeneous networks, leveraging their shared information is crucial for enhancing statistical efficiency, whereas accurately characterizing unique features is equally important. The coexistence of shared and unique patterns calls for analysis capable of disentangling the intertwined structural features and still fully exploiting statistical efficiency. Recently, MacDonald et al., (2022) introduced a general class of latent space models for multiple networks that decompose latent vectors into shared and network-specific components. Despite flexibility of the models, fundamental statistical limits under them remain unclear. Specifically, a critical statistical question for jointly analyzing multiple networks is: Can pooling heterogeneous networks improve the efficiency of estimating latent vectors, and if so, how can such gains be effectively achieved?
@@
To address this, we develop a unified framework of efficient analysis under the latent space modeling of multiple heterogeneous networks. We propose a novel strategy that first hunts for the shared latent space across networks and then applies a refinement to achieve oracle efficiency. In this way, challenges from identifying shared space and achieving efficiency can be tackled separately, enhancing analytical flexibility. To theoretically quantify the efficiency gain, we develop novel techniques that disentangle the distinct oracle estimation error rates of shared and network-specific spaces through analyzing complex efficient score equations. Our proposed framework is versatile and encompasses a broad range of edge weight distributions.
@@
We summarize notations used below. Given two sequences of real numbers $\{g_{n}\}$ and $\{h_{n}\}$, $g_{n}=O(h_{n})$ and $g_{n}\lesssim h_{n}$ represent $|g_{n}|\leqslant c_{1}|h_{n}|$ for a constant $c_{1}$, $g_{n}\ll h_{n}$ and $g_{n}=o(h_{n})$ mean $\lim_{n\to\infty}g_{n}/h_{n}=0$, and $g_{n}\gg h_{n}$ means $\lim_{n\to\infty}h_{n}/g_{n}=0$. For a sequence of random variables $\{X_{n}\}$, write $X_{n}=O_{p}(g_{n})$ if for any $\epsilon>0$, there exists finite $M>0$ such that $\sup_{n}\Pr(|X_{n}/g_{n}|>M)<\epsilon$; write $X_{n}=o_{p}(g_{n})$ if for any $\epsilon>0$, $\lim_{n\to\infty}\Pr(|X_{n}/g_{n}|>\epsilon)=0$. For $x=(x_{i})_{i=1}^{n}$ and $y=(y_{i})_{i=1}^{n}\in\mathbb{R}^{n}$, $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$, $\|x\|_{2}=\sqrt{\langle x,x\rangle}$, and $\|x\|_{\infty}=\max_{1\leqslant i\leqslant n}|x_{i}|$. For $X=(x_{ij})_{1\leqslant i\leqslant n,1\leqslant j\leqslant m}\in\mathbb{R}^{n\times m}$, $\|X\|_{\mathrm{F}}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}x_{ij}^{2}}$, $\|X\|_{\operatorname{op}}=\sup_{\|v\|_{2}=1}\|Xv\|_{2}$, and $\|X\|_{2\to\infty}=\sup_{\|v\|_{2}=1}\|Xv\|_{\infty}$. Let $\mathrm{I}_{k}$ denote $k\times k$ identity matrix, and $\mathcal{O}(k)=\left\{Q\in\mathbb{R}^{k\times k}:QQ^{\top}=\mathrm{I}_{k}\right\}$ represent the set of $k\times k$ orthogonal matrices.
@@ 2.1 Latent space model with shared factors
This paper focuses on multiple networks where each network is observed on a common set of $n$ codes without inter-layer connections. Suppose we observe $T$ undirected networks on a common set of $n$ nodes, represented by $n\times n$ adjacency matrices $\{\mathbf{A}_{t}=(A_{t,ij})_{1\leqslant i,j\leqslant n}:t=1,\ldots,T\}$. For weighted networks, $A_{t,ij}$ is the weight value, which could be continuous or a count number. For unweighted networks, $A_{t,ij}$ is binary indicating that the connection between nodes $i$ and $j$ exists or not.
@@
To model interactions between nodes, we adopt the popular latent space modeling framework (Hoff et al.,, 2002). Assume there exist latent vectors $y_{t,i}\in\mathbb{R}^{d_{t}}$ for $i=1,\ldots,n$ and $t=1,\ldots,T$ such that conditioning on $y_{t,i}$’s, network edges $A_{t,ij}$ are independently generated. To flexibly model different types of weight values, we allow $A_{t,ij}$ to be generated from a generic parametric distribution $p(\,\cdot\mid\theta\,)$ characterized by a parameter $\theta$, i.e.,
$A_{t,ij}=A_{t,ji}\ \sim\ p(\,\cdot\mid\Theta_{t,ij}\,),\quad\quad 1\leqslant i\leqslant j\leqslant n,\ 1\leqslant t\leqslant T,$ (1)
independently, where $\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle$ models the interaction effect between nodes $i$ and $j$ similarly to the popular low-rank models in the literature (Hoff et al.,, 2002; Ma et al.,, 2020). Examples of $p(\,\cdot\mid\theta\,)$ include Bernoulli, Gaussian, and Poisson distributions, commonly used to model edge weights that are binary, continuous, and counts, respectively. Model (1) posits a common $p(\,\cdot\mid\theta\,)$ across networks for notational simplicity. Our proposed analysis framework can be readily generalized when networks follow different types of distributions.
@@
For each node $i$, we assume that the latent vectors $\{y_{t,i}:t=1,\ldots,T\}$ may include shared and heterogeneous components over $T$ networks following MacDonald et al., (2022). Specifically, $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}\in\mathbb{R}^{k+k_{t}}$, where $k+k_{t}=d_{t}$, $z_{i}\in\mathbb{R}^{k}$ represents a shared component over $T$ networks, and $w_{t,i}\in\mathbb{R}^{k_{t}}$ represents a heterogeneous component that may differ across $1\leqslant t\leqslant T$. Then $\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle=\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle.$
@@
Let $\Theta_{t}=(\Theta_{t,ij})_{1\leqslant i,j\leqslant n}\in\mathbb{R}^{n\times n}$ denote the parameter matrix for the $t$-th network. We define $Y_{t}=[Z,\,W_{t}]\in\mathbb{R}^{n\times d_{t}},$ $Z=\left[z_{1},\ldots,z_{n}\right]^{\top}\in\mathbb{R}^{n\times k},$ and $W_{t}=\left[w_{t,1},\ldots,w_{t,n}\right]^{\top}\in\mathbb{R}^{n\times k_{t}}.$ Then our model has an equivalent low-rank matrix representation after link transformation
$\mu^{-1}\{\mathbb{E}(\mathbf{A}_{t})\}=\Theta_{t}=Y_{t}Y_{t}^{\top}=ZZ^{\top}+W_{t}W_{t}^{\top},\hskip 15.00002pt1\leqslant t\leqslant T,$ (2)
where $\mu(\cdot)$ denotes the unique invertible function linking $\theta$ and variable expectation under $p(\,\cdot\mid\theta\,)$, and it operates entrywisely on a matrix argument in (2).
Remark 1
The core model (1) is flexible and can be readily adapted to accommodate diverse network properties. For example, to allow flexible node degrees, an overall degree factor varying with respect to $(n,T)$ can be added. Also, (1) can be augmented with an extra probability component through mixture modeling, enabling features such as link missingness or disconnections in weighted networks. Our developments can be similarly established under these generalized models, exemplifying the broader applicability of our framework. Detailed results are provided in Sections J–L of the Supplementary Material.
@@ 2.2 Identifiability
By (2), the parameters $Z$ and $W_{t}$ can only be identified up to orthogonal transformation, i.e., $ZZ^{\top}=ZQQ^{\top}Z^{\top}$ and $W_{t}W_{t}^{\top}=W_{t}Q_{t}Q_{t}^{\top}W_{t}^{\top}$ for any $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$. We say that parameters $Z$ and $W_{t}$ in the model (2) are identifiable up to orthogonal group transformation if, for any two sets of parameters $\{Z,W_{1},\ldots,W_{T}\}$ and $\{Z^{\prime},W_{1}^{\prime},\ldots,W_{T}^{\prime}\}$ yielding the same model (2), there exist $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$ such that $Z=Z^{\prime}Q$ and $W_{t}=W^{\prime}_{t}Q_{t}$ for $1\leqslant t\leqslant T$. We next establish a sufficient condition for the identifiability of the model (2).
@@
    there exist $1\leqslant t<s\leqslant T$ such that the columns of $Z,W_{t},W_{s}]$ are linearly independent.
@@
Proposition 1 is adapted from MacDonald et al., (2022) but relaxes the connectivity assumption of the inter-layer graphs. Given (i) and (ii) in Proposition 1, the column span of $Z$ equals the intersection of all column spans of $Y_{t}$ across $1\leqslant t\leqslant T$.
@@ 2.3 Efficient analysis of latent spaces and challenges
For estimating the shared factor $Z$ under (2), we expect a similar semiparametric oracle error rate. As an illustration, consider an oracle scenario when heterogeneous (nuisance) parameters $w_{t,i}$’s are known. In this case, each shared (target) parameter $z_{i}$ is measured by $nT$ independent edges $\{A_{t,ij}:j=1,\ldots,n;t=1,\ldots,T\}$. The oracle squared estimation error rate of each latent vector $z_{i}$ is expected to be of the order of $O_{p}(k/(nT))$, the ratio between the parameter number and the effective sample size (Portnoy,, 1988). Thus the aggregated oracle estimation error of $n$ latent vectors $Z$ is expected to be $O_{p}(n\times k/(nT))=O_{p}(1/T)$ under fixed $k$. Theoretically, achieving this oracle error rate requires approximating Cramér-Rao lower bound for high-dimensional $Z$; more discussions on the notion of efficiency are given in Remark D.4 of the Supplementary Material. Nevertheless, to establish the result, technical challenges similar to those in He et al., (2025) persist.
@@
    Oracle error rates for estimating target and nuisance parameters are different, but their errors are entangled in the analysis. Intuitively, each nuisance $w_{t,i}$ is measured by $n$ independent edges: $\left\{A_{t,ij}:j=1,\ldots,n\right\}$, so the oracle squared error rate of each $w_{t,i}$ is expected to be $O_{p}(1/n)$, larger than $O_{p}(1/(nT))$ of each $z_{i}$. In theoretical analysis, it is difficult to separate the error of $z_{i}$ from that of $w_{t,i}$ due to several reasons. One is that the individual components $w_{t,i}$’s encode both node-specific heterogeneity over $i$ and network-specific heterogeneity over $t$. As a result, classical partial likelihood approach (Andersen and Gill,, 1982) cannot directly remove the heterogeneous factors. Moreover, when the link function in (2) is non-linear, such as the logistic or exponential function, semiparametric oracle rates for $Z$ cannot be easily obtained by spectral-based analyses (Arroyo et al.,, 2021) or analysis under Gaussian distribution (MacDonald et al.,, 2022).
@@
    Target parameters are unidentifiable. As argued in Section 2.2, the shared (target) latent space $Z$ is unidentifiable in the Euclidean space. Actually, the intrinsic space of $Z$ is a quotient set with the equivalence relation up to the orthogonal group transformation. To evaluate the discrepancy between a target matrix $Z^{\star}$ and an estimate $\hat{Z}$, a natural distance is $\mathrm{dist}(\hat{Z},Z^{\star})=\min_{Q\in\mathcal{O}(k)}\|\hat{Z}-Z^{\star}Q\|_{\mathrm{F}}$, instead of the Euclidean distance.
@@
    Both target and nuisance parameters are high-dimensional. Specifically, under (2), the dimension of $Z$ increases with respect to the network size $n$, and the total number of parameters in $W_{1},\ldots,W_{T}]$ grows with respect to both $n$ and $T$.
On the other hand, our model (2) posits unique challenges and properties, making existing analyses inapplicable and necessitating new developments of methodology and theory. We summarize critical distinctions into four key aspects.
@@
    Intricate factor interrelationships. The relationship between the shared (target) space $Z$ and individual (nuisance) spaces $W_{t}$’s exhibits unprecedented complexity. In He et al., (2025), the target and nuisance parameters are latent factors and baseline degrees, respectively, which characterize intrinsically orthogonal information that can enhance theoretical analysis. In contrast, factor interrelationships in (2) defy simple characterizations, which will be further elaborated in Remark 4. This complexity greatly amplifies the challenge of separating estimation errors across different $1+T$ factor groups.